ADNAN MUHAMMED

Grade 12,

sir pls give solution Q. find the value of x for which 2 2x+3 – 65(2 x -1)=57 -3 3 9 -9

sir pls give solution
Q. find the value of x for which 22x+3 – 65(2x-1)=57
  1. -3
  2. 3
  3. 9
  4. -9

Grade:12th pass

1 Answers

Vikas TU
7 years ago
-1)=57x – 65(22x+3Write:  2
as  (2x)^2.8 – 65(2x-1)=57
=> and assume 2^x = t
=> 8t^2 – 65t = 57 – 65
=> 8t^2 – 65t = -9
solve the quadratic.